Determine whether the point (8, 6) is a solution of the system
11x – 8y = 40
-6x + 1y = -42
by substituting the point into each equation.

Answer by rfer(16322)

11*8-8*6=40
88-40=48
doesn't work
--------------
-6*8+1*6=-42
-48+6=-42
Works on this one but not the first